# Tip growth in a strongly concentrated aggregation model follows local geodesics

### v2

@@ Abstract
We analyse the aggregate Loewner evolution (ALE), introduced in 2018 by Sola, Turner and Viklund to generalise versions of diffusion limited aggregation (DLA) in the plane using complex analysis. They showed convergence of the ALE for certain parameters to a single growing slit. Started from a non-trivial initial configuration of $k$ needles and the same parameters, we show that the small-particle scaling limit of ALE is the Laplacian path model, introduced by Carleson and Makarov in 2002, in which the tips grow along geodesics towards $\infty$.
The proof of this result involves analysis of Loewner’s equation near its singular points, and we extend usual martingale methods to this setting. Most conformal growth models introduce an extra regularisation factor to deal with the singularities in Loewner’s equation at the sharp tips and right-angle bases of slit particles. As an intermediate step we prove a limit result for a model with no such regularisation factor, developing methods which should prove useful in analysing other weakly-regularised models with non-trivial limits.
@@ 1 Introduction
We will prove that the ALE model of particle aggregation converges, in the limit as individual particles become small, to the LPM. This is an extension of the result of [14], that ALE started from a trivial initial configuration converges to an LPM with one arm, i.e. to a straight line. Our methods involve an explicit analysis of the tips of particles, where usual martingale methods are difficult to apply due to singularities in Loewner’s equation which describes both the ALE and LPM.
Figure 1: Simulations of the aggregate Loewner evolution started from a non-trivial configuration of arms. The initial cluster is drawn with thin black lines, and the attached particles are coloured by arrival time (the lighter particles arrived earlier). In the top-left diagram $\sigma$ is large enough that some particles are not attached at the tips of the existing slits and we see “branching” behaviour which does not occur in the limiting regime. In the three diagrams on the left one slit seems to grow much slower than the rest, agreeing with the observation in [1] that there is competition between the arms in the Laplacian path model, with only a certain number (depending on $\eta$) likely to survive (i.e. growth to a length proportional to the diameter of the entire cluster) as $T\to\infty$. The code used for the simulations is available at https://github.com/frankiehiggs/ALE-from-slits.
@@ 1.1 Loewner’s equation
Loewner’s equation encodes families of growing sets by measures on the cylinder $\mathbb{T}\times[0,T]$. Continuity properties of the encoding allow us to show the sets are close if the corresponding measures are close.
@@
Let $\mathbb{D}=\{z\in\mathbb{C}:|z|<1\}$ be the open unit disc, and $\Delta=\mathbb{C}_{\infty}\setminus\overline{\mathbb{D}}$ the open exterior disc, the complement of the closed unit disc in the Riemann sphere.
@@
[…]
@@
Let $\xi:[0,T]\to\mathbb{R}$ be a càdlàg function. Then there is a unique solution $\varphi:[0,T]\times\Delta\to\Delta$ to Loewner’s equation,
@@
Given a family of probability measures $(\mu_{t})_{t\geq 0}$, subject to measurability conditions on $t\mapsto\mu_{t}$ there is a unique solution to Loewner’s equation
@@ 1.2 Particle aggregation models
Any cluster constructed this way can be encoded by Loewner’s equation. For $t\geq 0$ let $n_{t}=\inf\{n\geq 1:\sum_{k=1}^{n}c_{k}>t\}$ (if $c_{n}=\mathbf{c}$ for all $n$, then $n_{t}=\left\lfloor t/\mathbf{c}\right\rfloor+1$). Let $\xi_{t}:=\theta_{n_{t}}$, and let $(\varphi_{t})_{t\geq 0}$ be the solution to (1.1) with driving measure $(\xi_{t})_{t\geq 0}$. Then for all $n$, $\varphi_{c(K_{n})}=\Phi_{n}$.
@@ 1.3 Aggregate Loewner evolution
If $(\theta_{n})_{n\geq 1}$ and $(c_{n})_{n\geq 1}$ are stochastic processes, then $(K_{n})_{n\geq 1}$ as constructed in Section 1.2 is a stochastic process in the space of compact subsets of $\mathbb{C}$. The aggregate Loewner evolution ALE($\alpha,\eta,\sigma$), introduced in [14] is one such process.
@@
We have introduced three parameters: $\alpha\in\mathbb{R}$, $\eta\in\mathbb{R}$, and $\sigma>0$. In this paper we will work with $\alpha=0$, which means that all particles have the same capacity.
@@ 1.4 Laplacian path model
where $e^{i\phi^{j}_{t}}$ is the preimage of $a^{j}_{t}$ under $\Phi_{t}^{\mathrm{LPM}}$, and $p^{j}_{t}=\frac{|(\Phi_{t}^{\mathrm{LPM}})^{\prime\prime}(e^{i\phi^{j}_{t}})|^{-\eta}}{Z_{t}}$, and $Z_{t}=\sum_{j=1}^{k}|(\Phi_{t}^{\mathrm{LPM}})^{\prime\prime}(e^{i\phi^{j}_{t}})|^{-\eta}$.
@@
    We could also derive $\mu_{t}^{\mathrm{LPM}}$ by removing the regularisation from the ALE attachment measure. For $\sigma>0$, let $\mu_{t}^{\sigma}$ be the regularised ALE measure on $\mathbb{T}$ corresponding to the LPM cluster at time $t$, i.e. the measure with density $h^{\sigma}(\theta)\propto|(\Phi_{t}^{\mathrm{LPM}})^{\prime}(e^{\sigma+i\theta})|^{-\eta}$. If $\eta>1$ then as $\sigma\to 0$ the density becomes concentrated around the zeroes of $(\Phi_{t}^{\mathrm{LPM}})^{\prime}$ at $\phi_{t}^{j}$ for $j=1,\dots,k$, and $|(\Phi_{t}^{\mathrm{LPM}})^{\prime}(e^{\sigma+i\phi^{j}_{t}})|\sim\sigma|(\Phi_{t}^{\mathrm{LPM}})^{\prime\prime}(e^{i\phi_{t}^{j}})|$. Hence $\mu_{t}^{\sigma}\to\mu_{t}^{\mathrm{LPM}}$ as $\sigma\to 0$ for a fixed time $t$.
  • The $\eta$ above corresponds to $\eta-1$ in the definitions of [1]; we have shifted it to match the $\eta$ used in the definition of the ALE in [14].
  • For any configuration of $k$ slits, $K_{0}$ is a “gearlike domain” as defined in [11], and so Schwarz–Christoffel methods can be used to compute an expression for $\Phi_{0}^{\mathrm{LPM}}$. To compute the expression, the preimages of all the tips and bases of the slits must be determined, which can be done numerically using methods in [11]. We used this to compute the maps for arbitrary slit configurations in our simulation of the ALE shown in Figure 1.
@@
Although little progress has been made since 2002 on analysing the geodesic LPM, various modifications have been studied and applied. In [5], Gubiec and Szymczak apply a similar construction to model finger growth in the half-plane.
The LPM and similar models have also been used to model the development of cracks in materials and formation of systems of rivers and streams. In particular, the chordal geodesic LPM with two needles at the tip of a half-line can be used to model the bifurcation of a stream. Carleson and Makarov proved that the angle between the two resulting streams in the chordal LPM must be $52\pi/$ 5 (also predicted by other authors using conformal mappings [6] [3]), which agrees with the average angle of $72^{\circ}$ measured in a Florida stream system [2] [16].
@@ 1.5 Main result
Let $(\Phi^{\mathrm{LPM}}_{t})_{t\in[0,T]}$ be the map for the LPM started from the same initial conditions, and let $\mu^{\mathrm{LPM}}_{t}$ be the driving measure $\mu^{\mathrm{LPM}}_{t}=\sum_{j}\bar{p}^{j}_{t}\delta_{\bar{\phi}^{j}_{t}}$.
@@
We will then introduce the multinomial model to replace the regularised derivative in the ALE with the second derivative on the boundary. This is quite simple. The most difficult and involved step is showing that the multinomial model, which is a discrete-time random growth model, converges to the continuous-time deterministic LPM. We essentially prove a strong law of large numbers at each tip: if we have attached a large number of slits by time $t$ and the probability of attaching at slit $j$ is $p$, then this is similar to a solution to Loewner’s equation at time $t$ with a driving function that always gives weight $p$ to the tip of slit $j$. It is complicated by the fact that the probabilities and weights both change over time, as do the preimages of the tips. Another important technical detail is that Loewner’s equation is singular at the tips.
@@ 2 Analytic tools
$f(z)=\frac{e^{\mathbf{c}}}{2w}(w+1)^{2}\left(1+\sqrt{\frac{w^{2}+2(1-2e^{-\mathbf{c}})w+1}{w^{2}+1}}\right)-1$ (2.1)
@@
Near $e^{\pm i\beta}$, $f$ has a non-linear effect on distances, quantified in [8, Lemma 12]:
@@
[…]
One useful tool for analysing Loewner’s equation is the backward equation. See [10] for more details.
@@
If the driving function $\xi$ corresponds to an aggregation process, so $\xi_{t}=\theta_{\left\lfloor t/\mathbf{c}\right\rfloor+1}$, then solving Loewner’s (forward) equation gives $\varphi_{N\mathbf{c}}=f_{1}\circ\dots\circ f_{N}$. Understanding $\varphi_{N\mathbf{c}}(z)$, $\varphi_{N\mathbf{c}}^{\prime}(z)$, etc. requires analysing $f_{N}(z)$, $f_{N-1}(f_{N}(z))$, etc.. For example, $\varphi_{N\mathbf{c}}^{\prime}(z)=\prod_{n=1}^{N}f_{n}^{\prime}((f_{n+1}\circ\dots\circ f_{N})(z))$, but the location of $(f_{n+1}\circ\dots\circ f_{N})(z)$ is difficult to understand using the forward equation. However, if we solve the backward equation (2.3), we have $u_{n\mathbf{c}}=f_{N-n+1}\circ\dots\circ f_{N}$ for all $n\leq N$, so we could write $\varphi_{N\mathbf{c}}^{\prime}(z)=\prod_{n=1}^{N}f_{n}^{\prime}(u_{(N-n)\mathbf{c}}(z))$.
@@
The estimate (2.4) follows by expanding (2.2). When $|\theta|>\beta$, [14, equation (37)] says that
@@
(where we interpret the left hand side as being equal to 00 if $z=z_{0}$) and
@@ 3 Removing the regularisation
In the following subsections we reduce the ALE, which has an attachment density supported on $\partial\mathbb{T}$, to the “multinomial model,” which, like the LPM, grows only at the tips, and the probability of attaching to a tip at time $t$ is determined by the second derivative of the map corresponding to the cluster.
@@ 3.1 ALE to auxiliary model
Figure 2: One step in the construction of the auxiliary model. The large red dot represents the image of $\exp(i\theta_{6}^{*})$ under each map. Comparing $K_{5}$ and $K_{6}$, we can see that both have three curves but the attachment points have been rotated from one cluster to the next. For this symmetric initial configuration and small capacity, the curves resemble slits, but this would not be true in general.
@@
To show that the ALE and auxiliary models are close, we will show that with high probability they land near the same slit, and the ALE lands very near the tip.
@@
The ALE and auxiliary models will have a common weak limit if with high probability $\tau_{D}\wedge\tau_{\mathrm{coupling}}>T/\mathbf{c}$. This event means that every particle in the ALE model has been attached within $D$ of a “main” tip, and every particle in the auxiliary model has chosen the same slit to attach to as the ALE model.
This is the main result of this section:
@@
provided $\sigma<\mathbf{c}^{\gamma}$ for some $\gamma\geq\frac{2(\eta+2)}{\eta-1}\vee\frac{5\eta+10}{2\eta}\vee 8$.
@@
The three terms determining $\gamma$ come from three separate requirements in the proof of this section’s result: we require $\sigma<\mathbf{c}^{8}$ so that the derivatives of each model look similar near the “main” tips, as in the following lemma. We require that $\sigma<\mathbf{c}^{\frac{5\eta+10}{2\eta}}$ so that “old” particles do not contribute in the ALE model, and we require that $\sigma<\mathbf{c}^{\frac{2(\eta+2)}{\eta-1}}$ so that the measures are concentrated very tightly around the tip of each particle, with each particle attached within distance $D$ of a main tip.
@@
It would be possible to weaken the conditions on $\gamma$ if we replaced the $A\mathbf{c}^{1/2}$ with a weaker $o(1)$ upper bound in Theorem 3.1 (which is all we require to prove Theorem 1.1) but we have left it in its current form to avoid complicating the proofs further.
@@
If $\sigma<\mathbf{c}^{\gamma}$ where $\gamma\geq\frac{2(\eta+2)}{\eta-1}\vee\frac{5\eta+10}{2\eta}\vee 8$, then this implies $\mathbb{P}[n+1=\tau_{D}\,|\,n<\tau_{D}]\leq A\mathbf{c}^{2}$ and so $\mathbb{P}[\tau_{D}\leq\left\lfloor T/\mathbf{c}\right\rfloor]\leq AT\mathbf{c}$.
@@
Now we have to prove Propositions 3.2 and 3.3. We begin with Proposition 3.3 as its proof is somewhat simpler.
@@
where for $0<\ell<N_{j}$, $\Psi^{j}_{\ell}=f^{\mathrm{ALE}}_{n_{j}(\ell)+1}\circ f^{\mathrm{ALE}}_{n_{j}(\ell)+2}\circ\dots\circ f^{\mathrm{ALE}}_{n_{j}(\ell+1)-1}=\left(\Phi^{\mathrm{ALE}}_{n_{j}(\ell)}\right)^{-1}\circ\Phi^{\mathrm{ALE}}_{n_{j}(\ell+1)-1}$, the map which attaches every particle landing at slits other than the $j$th between the $\ell$th and $(\ell+1)$th time a particle lands on the $j$th slit. The two other maps are $\Psi^{j}_{0}=\left(\Phi^{\mathrm{ALE}}_{0}\right)^{-1}\circ\Phi^{\mathrm{ALE}}_{n_{j}(1)-1}$ and $\Psi^{j}_{N_{j}}=\left(\Phi^{\mathrm{ALE}}_{n_{j}(N_{j})}\right)^{-1}\circ\Phi^{\mathrm{ALE}}_{n}$, which correspond to the particles before the first attachment and after the last attachment to the $j$th slit, respectively. Note that any of these $\Psi^{j}_{\ell}$ maps may be the identity map on $\Delta$ if $n_{j}(\ell+1)=n_{j}(\ell)+1$. The important common feature of all these $\Psi^{j}_{\ell}$ maps is that their derivatives have no poles or zeroes near the $j$th slit.
@@
We know $e^{i\phi^{j}_{n}}=\left(\Psi^{j}_{N_{j}}\right)^{-1}(\theta^{\mathrm{ALE}}_{n_{j}(N_{j})})$ and $e^{i\bar{\phi}^{j}_{n}}=\left(\overline{\Psi}^{j}_{N_{j}}\right)^{-1}(\hat{\theta}^{*}_{n_{j}(N_{j})}+\Delta_{n_{j}(N_{j})}^{n})$, and so applying Lemma 2.5 gives us, since $(n-n_{j}(N_{j}))\mathbf{c}\leq T$,
@@
It is clear from (2.2) that, to first order, the size of $|f^{\prime}(z)|$ depends on $|z-1|$. This means that to find the size of the two derivatives above, we are most interested in
@@
Firstly, a calculation using the explicit form of $f(z)$ in (2.1) shows that if $|z-1|<\mathbf{c}^{1/2}$ then $f(z)=1+d+O(|z-1|)$. The next estimate $z=1+O(|z-1|)$ is obvious. Putting this estimate into the denominator of the second fraction in (2.2), we get $|z-e^{\pm i\beta}|^{1/2}=|1-e^{i\beta}|^{1/2}(1+O\left(\frac{|z-1|}{\mathbf{c}^{1/2}}\right))$. Thus for $z$ close to 1 the behaviour of $|f^{\prime}(z)|$ mainly depends on $|z-1|$, and we can write

### v3

@@ Abstract
We analyse the aggregate Loewner evolution (ALE), introduced in 2018 by Sola, Turner and Viklund to generalise versions of diffusion limited aggregation (DLA) in the plane using complex analysis. They showed convergence of the ALE for certain parameters to a single growing slit. Started from a non-trivial initial configuration of $k$ needles and the same parameters, we show that the small-particle scaling limit of ALE is the Laplacian path model, introduced by Carleson and Makarov, in which the tips grow along geodesics towards $\infty$.
Our proof involves analysis of Loewner’s equation near its singular points, and we extend martingale methods to the backward equation, where what we have to control is non-adapted. Most conformal growth models introduce an extra regularisation factor to deal with the singularities in Loewner’s equation at the sharp tips and right-angle bases of slit particles. As an intermediate step we prove a limit result for a model with no such regularisation factor, developing methods which should prove useful in analysing other weakly-regularised models with non-trivial limits.
@@ 1 Introduction
We will prove that the ALE model of particle aggregation converges, in the limit as individual particles become small, to the LPM. This is an extension of the result of [1], that ALE started from a disc with no slits initially attached converges to an LPM with one arm, i.e. to a straight line. Our methods involve an explicit analysis of the tips of particles, where usual martingale methods are difficult to apply due to singularities in Loewner’s equation which describes both the ALE and LPM.
Figure 1: Simulations of the aggregate Loewner evolution started from a non-trivial configuration of arms. The initial cluster is drawn with thin black lines, and the attached particles are coloured by arrival time (the lighter particles arrived earlier). In the top-left diagram $\sigma$ is large enough that some particles are not attached at the tips of the existing slits and we see “branching” behaviour which does not occur in the limiting regime. In the three diagrams on the left one slit seems to grow much slower than the rest, agreeing with the observation in [2] that there is competition between the arms in the Laplacian path model, with only a certain number (depending on $\eta$) likely to survive (i.e. grow to a length proportional to the diameter of the entire cluster) as $T\to\infty$. The code used for the simulations is available at https://github.com/frankiehiggs/ALE-from-slits
2 Preliminaries
@@ 2.1 Loewner’s equation
Loewner’s equation encodes families of growing sets by measures on a cylinder. Continuity properties of the encoding allow us to show the sets are close if the corresponding measures are close.
@@
Let $\mathbb{D}=\{z\in\mathbb{C}:|z|<1\}$ be the open unit disc, and $\Delta=\mathbb{C}_{\infty}\setminus\overline{\mathbb{D}}$ the open exterior disc, the complement of the closed unit disc in the Riemann sphere. Let $\mathbb{T}=\partial\mathbb{D}$ denote their common boundary, the unit circle.
@@
If $K$ is a strict superset of $\overline{\mathbb{D}}$, then c(K)>0c(K)>0.111This follows from the fact that $c(K)=\int_{0}^{1}\log|f_{K}(e^{2\pi it})|\,\mathrm{d}t$. Another useful property of logarithmic capacity is additivity: if $f_{K}$ can be written $f_{K_{1}}\circ f_{K_{2}}$, then $c(K)=c(K_{1})+c(K_{2})$.
@@
Let $\xi:[0,T]\to\mathbb{R}$ be a càdlàg function [^2] A function $f:[0,T]\to\mathbb{R}$ is said to be càdlàg (right-continuous with left limits) if, for every $x\in[0,T]$ and every decreasing sequence $x_{n}\downarrow x$ and increasing sequence $y_{n}\uparrow x$, $f(x_{n})\to f(x)$ as $n\to\infty$ and $\lim_{n\to\infty}f(y_{n})$ exists (but is not necessarily equal to $f(x)$).. Then there is a unique solution $\varphi:[0,T]\times\Delta\to\Delta$ to Loewner’s equation,
@@
For the existence and uniqueness of the solution to (2.1), see [3, Theorem 3.4] (note that every càdlàg function is piecewise continuous, meeting the conditions in [3]).
@@
Given a family of probability measures on $\mathbb{T}$, $(\mu_{t})_{t\geq 0}$, subject to measurability conditions on $t\mapsto\mu_{t}$ there is a unique solution to Loewner’s equation
@@ 2.2 Particle aggregation models
Any cluster constructed this way can be encoded by Loewner’s equation. For $t\geq 0$ let $n_{t}=\inf\{n\geq 1:\sum_{k=1}^{n}c_{k}>t\}$ (if $c_{n}=\mathbf{c}$ for all $n$, then $n_{t}=\left\lfloor t/\mathbf{c}\right\rfloor+1$). [^3] If $\sum_{k=1}^{\infty}c_{k}<\infty$, then we can solve (2.1) up to a finite time. Let $\xi_{t}:=\theta_{n_{t}}$, and let $(\varphi_{t})_{t\geq 0}$ be the solution to (2.1) with driving measure $(\xi_{t})_{t\geq 0}$. Then for all $n$, $\varphi_{c(K_{n})}=\Phi_{n}$.
@@ 2.3 Aggregate Loewner evolution
If $(\theta_{n})_{n\geq 1}$ and $(c_{n})_{n\geq 1}$ are stochastic processes, then $(K_{n})_{n\geq 1}$ as constructed in Section 2.2 is a stochastic process in the space of compact subsets of $\mathbb{C}$. The aggregate Loewner evolution ALE($\alpha,\eta,\sigma$), introduced in [1], is one such process.
@@
We have introduced three parameters: $\alpha\in\mathbb{R}$, $\eta\in\mathbb{R}$, and $\sigma>0$. In this paper we will work with $\alpha=0$, which means that all particles have the same capacity. After the statement of Theorem 2.1 we will discuss the effect of changing $\alpha$.
@@ 2.4 Laplacian path model
where $e^{i\phi^{j}_{t}}$ is the preimage of $a^{j}_{t}$ under $\Phi_{t}^{\mathrm{LPM}}$, and
$p^{j}_{t}:=\frac{|(\Phi_{t}^{\mathrm{LPM}})^{\prime\prime}(e^{i\phi^{j}_{t}})|^{-\eta}}{Z_{t}},$ (2.4)
and
$Z_{t}:=\sum_{j=1}^{k}|(\Phi_{t}^{\mathrm{LPM}})^{\prime\prime}(e^{i\phi^{j}_{t}})|^{-\eta}.$ (2.5)
@@
We could also derive $\mu_{t}^{\mathrm{LPM}}$ by removing the regularisation from the ALE attachment measure. For $\sigma>0$, let $\mu_{t}^{\sigma}$ be the regularised ALE measure on $\mathbb{T}$ corresponding to the LPM cluster at time $t$, i.e. the measure with density $h^{\sigma}(\theta)\propto(\Phi_{t}^{\mathrm{LPM}})^{\prime}(e^{\sigma+i\theta})|^{-\eta}$. If $\eta>1$ then as $\sigma\to 0$ the density becomes concentrated around the zeroes of $(\Phi_{t}^{\mathrm{LPM}})^{\prime}$ at $\phi_{t}^{j}$ for $j=1,\dots,k$, and $|(\Phi_{t}^{\mathrm{LPM}})^{\prime}(e^{\sigma+i\phi^{j}_{t}})|\sim\sigma|(\Phi_{t}^{\mathrm{LPM}})^{\prime\prime}(e^{i\phi_{t}^{j}})|$. Hence $\mu_{t}^{\sigma}\to\mu_{t}^{\mathrm{LPM}}$ as $\sigma\to 0$ for a fixed time $t$.
Remark.
The $\eta$ in Definition 2.4.1 corresponds to $\eta-1$ in the definition given in [2]; we have shifted it to match the $\eta$ used in the definition of the ALE in [1].
Remark.
For any configuration of $k$ slits, $K_{0}$ is a “gearlike domain” as defined in [6], and so Schwarz–Christoffel methods can be used to compute an expression for $\Phi_{0}^{\mathrm{LPM}}$. To compute the expression, the preimages of all the tips and bases of the slits must be determined, which can be done numerically using methods from [6]. We used this to compute the maps for arbitrary slit configurations in our simulation of the ALE shown in Figure 1.
@@
Although little progress has been made since 2002 on analysing the geodesic LPM, various modifications have been studied and applied. In [8], Gubiec and Szymczak apply a similar construction to model finger growth in the half-plane. We hope that relating the LPM to the more active area of conformal growth models will revive interest in the model and many of the open problems stated in [2].
The LPM and similar models have also been used to model the development of cracks in materials and formation of systems of rivers and streams. In particular, the chordal geodesic LPM with two needles at the tip of a half-line can be used to model the bifurcation of a stream. Carleson and Makarov proved that the angle between the two resulting streams in the chordal LPM must be $2\pi/5$ (also predicted by other authors using conformal mappings [9, 10]), which agrees with the average angle of $72^{\circ}$ measured in a Florida stream system [11, 12].
@@ 2.5 Main result
Let $(\Phi^{\mathrm{LPM}}_{t})_{t\in[0,T]}$ be the map for the LPM started from the same initial conditions with the same parameter $\eta$, and let $\mu^{\mathrm{LPM}}_{t}$ be the driving measure $\mu^{\mathrm{LPM}}_{t}=\sum_{j}\bar{p}^{j}_{t}\delta_{\bar{\phi}^{j}_{t}}$.
@@
Remark.
Convergence of $K_{n}\to K$ in the Carathéodory topology is equivalent to convergence of the corresponding maps $f_{K_{n}}\to f_{K}$ on compact subsets of $\Delta$ [13, Chapter 3.6].
Remark.
In [1, Corollary 10] it is shown that if $\alpha>0$, the ALE started from the trivial initial cluster $\overline{\mathbb{D}}$ still converges to the union of $\overline{\mathbb{D}}$ with a single slit. We conjecture that a similar modification of our Theorem 2.1 should also be true:
Conjecture 2.2.
[…]
We expect that proving Conjecture 2.2 would be similar to the $\alpha=0$ case. The computations in Section 4 would become messier and the martingale arguments in Section 5.2 become a bit more difficult but fundamentally the same arguments should apply.
@@
We will then introduce the multinomial model to replace the regularised derivative in the ALE with the second derivative on the boundary. This is quite simple. The most difficult and involved step is showing that the multinomial model, which is a discrete-time random growth model, converges to the continuous-time deterministic LPM. We essentially prove a strong law of large numbers at each tip: if we have attached a large number of slits by time $t$ and the probability of attaching at slit $j$ is $p$, then this is similar to a solution to Loewner’s equation at time $t$ with a driving function that always gives weight $p$ to the tip of slit $j$. It is complicated by the fact that the probabilities and weights both change over time, as do the preimages of the tips. Analysing the fluctuation of the weights at each tip from the Laplacian path model’s (2.4) is the most technical part of Section 5.2, involving controlling strong feedback in the dynamics. Another important technical detail is that Loewner’s equation is singular at the tips.
@@ 3 Analytic tools
Remark.
Frequently throughout the paper, we will use a constant $A$ which changes from line to line, having the properties that $0<A<\infty$ and that $A$ depends only on the parameters which we are holding fixed, such as $T$, $K_{0}$ and $\eta$ in Theorem 2.1.
We will occasionally note explicitly on which parameters $A$ depends, but it is always independent of the parameter $\mathbf{c}$ which we allow to tend to 0, and of the number of particles.
@@
$f(z)=\frac{e^{\mathbf{c}}}{2w}(w+1)\left(w+1+\sqrt{w^{2}+2(1-2e^{-\mathbf{c}})w+1}\right)-1$ (3.1)
@@
Near $e^{\pm i\beta}$, $f$ has a non-linear effect on distances, quantified in [5, Lemma 12]:
@@
$\begin{aligned} |f(w)-1|=\ &2(e^{\mathbf{c}}-1)^{1/4}|w-e^{i\beta}|^{1/2}\\ &\times\left(1+O\left[\frac{|w-e^{i\beta}|}{\mathbf{c}^{1/2}}\vee\mathbf{c}^{1/4}|w-e^{i\beta}|^{1/2}\right]\right).\end{aligned}$
One useful tool for analysing Loewner’s equation is the backward equation. See [13, Chapter 4] [^5] The backward equation appears in [13] in the proof of Theorem 4.13..
@@
If the driving function $\xi$ corresponds to an aggregation process, so $\xi_{t}=\theta_{\left\lfloor t/\mathbf{c}\right\rfloor+1}$, then solving Loewner’s (forward) equation gives $\varphi_{N\mathbf{c}}=f_{1}\circ\dots\circ f_{N}$. Understanding $\varphi_{N\mathbf{c}}(z)$, $\varphi_{N\mathbf{c}}^{\prime}(z)$, etc. requires analysing $f_{N}(z)$, $f_{N-1}(f_{N}(z))$, etc. For example, $\varphi_{N\mathbf{c}}^{\prime}(z)=\prod_{n=1}^{N}f_{n}^{\prime}((f_{n+1}\circ\dots\circ f_{N})(z))$, but the location of $(f_{n+1}\circ\dots\circ f_{N})(z)$ is difficult to understand using the forward equation. However, if we solve the backward equation (3.3), we have $u_{n\mathbf{c}}=f_{N-n+1}\circ\dots\circ f_{N}$ for all $n\leq N$, so we could write $\varphi_{N\mathbf{c}}^{\prime}(z)=\prod_{n=1}^{N}f_{n}^{\prime}(u_{(N-n)\mathbf{c}}(z))$.
@@
The estimate (3.4) follows by expanding (3.2). When $|\theta|>\beta$, [1, equation (37)] says that
@@
(where we interpret the left hand side as being equal to 0 if $z=z_{0}$) and
@@ 4 Removing the regularisation
In this section we reduce the ALE, which has an attachment density supported on $\mathbb{T}$, to the “multinomial model,” which, like the LPM, grows only at the tips, and the probability of attaching to a tip at time $t$ is determined by the second derivative of the map corresponding to the cluster.
The proofs in this section mostly involve careful calculations to estimate $|\Phi_{n}^{\prime}|$ around relevant points.
@@ 4.1 ALE to auxiliary model
Figure 2: One step in the construction of the auxiliary model. The large dot at one tip represents the image of $\exp(i\theta_{6}^{*})$ under each map. Comparing $K_{5}$ and $K_{6}$, we can see that both have three curves but the attachment points have been rotated from one cluster to the next.
@@
To show that the ALE and auxiliary models are close, we will couple the models so that at each step the probability $j_{n}=j_{n}^{*}$ is close to 1, and show that with high probability the ALE lands very near a tip.
@@
The ALE and auxiliary models will have a common weak limit if with high probability $\tau_{D}\wedge\tau_{\mathrm{coupling}}>T/\mathbf{c}$. This event means that every particle in the ALE model has been attached within $D$ of a “main” tip, and every particle in the auxiliary model has chosen the same slit to attach to as the ALE model. An example of the coupling of these two models is shown in Figure 3.
@@
provided $\sigma<\mathbf{c}^{\gamma}$ for $\gamma=\frac{2(\eta+2)}{\eta-1}\vee\frac{5\eta+10}{2\eta}\vee 8$.
Figure 3: Coupled ALE (left) and auxiliary (right) models at a time $n<\tau_{D}\wedge\tau_{\mathrm{coupling}}$ started from a three-arm configuration. Note the slight rotation of three initial arms in the auxiliary model, and that the curves are repelled from each other.
@@
The three terms determining $\gamma$ come from three separate requirements in the proof of this section’s result: we require $\sigma<\mathbf{c}^{8}$ so that the derivatives of each model look similar near the “main” tips, as in Proposition 4.3. We require that $\sigma<\mathbf{c}^{\frac{5\eta+10}{2\eta}}$ so that “old” particles do not contribute in the ALE model, and we require that $\sigma<\mathbf{c}^{\frac{2(\eta+2)}{\eta-1}}$ so that the measures are concentrated very tightly around the tip of each particle, with each particle attached within distance $D$ of a main tip.
All three terms and their maximum are plotted in Figure 4.
@@
It would be possible to use a slightly smaller $\gamma$ if we replaced the $A\mathbf{c}^{1/2}$ with a weaker $o(1)$ upper bound in Theorem 4.1 (which is all we require to prove Theorem 2.1) but we have left it in its current form to avoid complicating the proofs further.
Figure 4: A plot of the exponent $\gamma$ used to control the regularisation in Theorem 4.1 as a function of the parameter $\eta$. Each of the three terms determining $\gamma$ is plotted separately, with their maximum shown as a solid black curve.
@@
If $\sigma<\mathbf{c}^{\gamma}$ where $\gamma=\frac{2(\eta+2)}{\eta-1}\vee\frac{5\eta+10}{2\eta}\vee 8$, then this implies $\mathbb{P}[n+1=\tau_{D}\,|\,n<\tau_{D}]\leq A\mathbf{c}^{2}$ and so $\mathbb{P}[\tau_{D}\leq\left\lfloor T/\mathbf{c}\right\rfloor]\leq AT\mathbf{c}$.
@@
Now we have to prove Propositions 4.2 and 4.3. We begin with Proposition 4.3 as its proof is somewhat simpler.
@@
[…]
@@
[…]
@@
It is clear from (3.2) that, to first order, the size of $|f^{\prime}(z)|$ depends on |z1||z-1|. This means that to find the size of the two derivatives above, we are most interested in
@@
Firstly, a calculation using the explicit form of $f(z)$ in (3.1) shows that if $|z-1|<\mathbf{c}^{1/2}$ then f(z)=f(1)+O(|z1|)=1+d+O(|z1|)f(z)=f(1)+O(|z-1|)=1+d+O(|z-1|). The next estimate z=1+O(|z1|)z=1+O(|z-1|) is obvious. Putting this estimate into the denominator of the second fraction in (3.2), we get $|z-e^{\pm i\beta}|^{1/2}=|1-e^{i\beta}|^{1/2}(1+O\left(\frac{|z-1|}{\mathbf{c}^{1/2}}\right))$. Thus for $z$ close to 1 the behaviour of $|f^{\prime}(z)|$ mainly depends on |z1||z-1|, and we can write